OpenStudy (angel_kitty12):

Can someone help me check this equation and make sure there are no extraneous solutions?

OpenStudy (angel_kitty12):

\[2\sqrt[3]{10-3(\frac{ 78 }{ 23})}=\sqrt[3]{2-(\frac{ 78 }{ 23})}\]

OpenStudy (angel_kitty12):

never mind. I figured it out. >:)